OpenStudy (anonymous):

Write an equation of the line that passes through the pair of points. (2,3), (4,-2)? a.y=- 5/2x-1/8 b.y=- 5/2 x-8 c.y= - 5/2x+8 d.y=5/2x+8

OpenStudy (anonymous):

Find slope first..

OpenStudy (anonymous):

i'm so confused:( @waterineyes

OpenStudy (anonymous):

Slope is given by: \[\huge \color{blue}{Slope = \frac{y_2-y_1}{x_2 - x_1}}\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Can you find slope now or not?? Just plug in the values, \[(x_1,y_1) = (2,3)\] \[(x_2,y_2) = (4,-2)\]

OpenStudy (anonymous):

I have to leave now... So, I am solving this for you: \[m = \frac{-2-3}{4-2} = \frac{-5}{2}\] The equation with one point and one slope is given by: \[\large \color{green}{y - y_1 = m(x - x_1)}\] Putting the first point \((2,3)\) in this: \[y - 3 = m(x - 2) \implies y - 3 = \frac{-5}{2}(x - 2)\] \[y - 3= \frac{-5}{2}x + 5 \implies y = \frac{-5}{2}x + 8\]